### v1

# Iband Measurements in a Sample of Hosts of 9 SN Ia Match HST Cepheids

Abstract
The Hubble Tension, a discrepancy between direct and indirect measurements of the Hubble constant (), has persisted for a decade and motivated intense scrutiny of the paths used to infer . Comparing independently-derived distances for a set of galaxies with different standard candles, such as the tip of the red giant branch (TRGB) and Cepheid variables, can test for systematics in the middle rung of the distance ladder. The band is the preferred filter for measuring the TRGB due to constancy with color, a result of low sensitivity to population differences in age and metallicity supported by stellar models. We use James Webb Space Telescope (JWST) observations with the maser host NGC 4258 as our geometric anchor to measure -band (F090W vs F090WF150W) TRGB distances to 7 hosts of 9 Type Ia supernovae (SNe Ia) within 27 Mpc: NGC 1448, NGC 1559, NGC 2525, NGC 3370, NGC 3447, NGC 5584, and NGC 5643. We compare these with Hubble Space Telescope (HST) Cepheid-based relative distance moduli for the same galaxies and anchor. We find no evidence of a difference between their weighted means, 0.04 (stat) (sys) mag. We produce fourteen variants of the TRGB analysis, altering the smoothing level and color range used to measure the tips to explore their impact. For some hosts, this changes the identification of the strongest peak, but this causes little change to the sample mean difference producing a full range of 0.01 to 0.03 mag, all consistent at 1 with no difference. The result matches past comparisons of -band TRGB and Cepheids when both use HST. SNe and anchor samples observed with JWST are too small to yield a measure of that is competitive with the HST sample of 42 SNe Ia and 4 anchors; however, they already provide a vital systematic crosscheck to HST measurements of the distance ladder.
1 Introduction
The ‘Hubble Tension’ refers to a 56 discrepancy between direct measurements and cosmological inference of the present expansion rate of the Universe, parameterized as the Hubble constant (). Currently, the “gold standard” method to directly measure utilizes a three-rung ladder to obtain distances to galaxies in the Hubble flow. This method uses geometric distances derived from parallaxes (such as those from Gaia Collaboration et al. 2021), water masers (Reid et al., 2019), and detached eclipsing binaries (Pietrzyński et al., 2019; Graczyk et al., 2020) to calibrate the luminosities of primary distance indicators such as Cepheid variables (first rung). This calibration is in turn used to measure distances to galaxies that host both Cepheids and Type Ia supernovae to calibrate the luminosities of the latter (second rung), which are then finally used to measure distances to galaxies in the Hubble flow to measure (third rung). The Supernovae, , for the Equation of State (SH0ES) team, for instance, has used this method to measure = 73.17 0.86 km s-1 Mpc-1 (Riess et al., 2022; Breuval et al., 2024). The cosmological inference uses the cosmic microwave background together with a cosmological model such as CDM to infer the present expansion rate. Planck Collaboration et al. (2020) used this approach to derive 67.36 0.54 km s-1 Mpc-1, a difference from the value obtained by Breuval et al. (2024). Several independent teams and measurements using a wide range of distance indicators have also yielded measurements in tension with those from Planck. No precise direct measurement of has yielded a value that is lower than that from Planck (see compilations in Di Valentino et al. 2021; Verde et al. 2023). This tension suggests a possible need to revise the standard CDM model, though efforts continue to search for unidentified systematics in the measurements at any level111For an extensive list of studies that have studied systematics related to the tension, we refer the reader to https://djbrout.github.io/SH0ESrefs.html..
Independent distance measurements to the same galaxies using multiple standard candles offers one of the best ways to ensure systematics are well accounted for and understood. The tip of the red giant branch (TRGB) refers to the maximum luminosity reached by first-ascent red giant stars before transitioning onto the horizontal branch, and has been used as a standard candle to measure distances (e.g. Anand et al., 2019a, b; Beaton et al., 2019; Hoyt et al., 2021; Lee et al., 2021a; Anand et al., 2021a, b; Oakes et al., 2022; Tran et al., 2022; Lee et al., 2022; Madore & Freedman, 2024; Anand et al., 2024a, b), and the Hubble constant (e.g. Freedman et al., 2019; Kim et al., 2020; Freedman, 2021a; Blakeslee et al., 2021; Dhawan et al., 2022; Anand et al., 2022; Dhawan et al., 2023; Scolnic et al., 2023, see also review in Valentino & Brout 2024). The TRGB can thus be used as a means to crosscheck systematics of other standard candles, such as Cepheid variables.
The TRGB is commonly measured in the I band, or the Hubble Space Telescope (HST) F814W and James Webb Space Telescope (JWST) F090W equivalents, where the magnitude of the TRGB exhibits the smallest metallicity (and hence color) dependence (Rizzi et al., 2007; Jang & Lee, 2017a). While it is possible to measure the TRGB in the NIR to take advantage of brighter magnitudes, doing so requires careful calibration of the TRGB’s color dependence (Bellazzini et al., 2004; Dalcanton et al., 2012; Wu et al., 2014; Madore et al., 2018; Hoyt et al., 2018; McQuinn et al., 2019; Newman et al., 2024a, b; Hoyt et al., 2024) where it rises to one magnitude per magnitude of measured color, introducing additional systematics. In this study, we focus on measuring the TRGB in the JWST F090W filter, which is most similar to I and F814W used with HST.
The launch of JWST has opened up the possibility of performing extensive crosschecks of Cepheids, TRGB, and other standard candles in unprecedented resolution. In particular, the dual-module configuration of the JWST NIRCam instrument allows for at least three standard candles (Cepheids, TRGB, and Carbon stars) to be simultaneously observed with one set of observations see for instance, JWST programs GO-1685 (Riess et al., 2021), GO-1995 (Freedman et al., 2021), and GO-2875 (Riess et al., 2023), among others. In a previous study, Anand et al. (2024a) (Paper I), we measured the TRGB with JWST in the maser host NGC 4258, which serves as one of four geometric calibrators of the distance ladder, and presented measurements in 2 SN Ia hosts (NGC 1559 and NGC 5584). The derived luminosity of the TRGB in JWST F090W in Paper 1 matched that independently determined by Newman et al. (2024a), also in F090W, to 0.01 mag, indicating a reliable foundation for distance measurements with JWST in SN Ia hosts. Here we study an expanded sample using JWST observations of 7 hosts of 9 SNe Ia from Cycle 1 program GO-1685 (Riess et al., 2021) and Cycle 2 GO-2785 (Riess et al., 2023) to measure TRGB-based distances to these galaxies and test their consistency with HST Cepheid-based distances to those same galaxies.
2 Data
We retrieve JWST F090W and F150W observations of NGC 1448, NGC 1559, NGC 5584, and NGC 5643 from GO-1685 (Riess et al., 2021) and NGC 2525, NGC 3370, and NGC 3447 from GO-2875 (Riess et al., 2023) using the Barbara A. Mikulski Archive for Space Telescopes (MAST)222https://mast.stsci.edu/portal/Mashup/Clients/Mast/Portal.html. We list observation details in Table 1 and show galaxy footprints in Fig. 1. For NGC 2525, we found that the images corresponding to the fourth (and final) dither for each filter were noticeably blurred. Upon further investigation, we found that the guide star had wandered onto some bad pixels in the Fine Guidance Sensor. We thus excluded these dithers from our analysis and generated our own stage 3 image from only the first three dithers for use of mutual image alignment of the underlying stage 2 images with DOLPHOT.
Galaxy | Program | Observation Number | Observation Date | Filters | Exposure Time [s] | pmap Version |
NGC 1448 | GO-1685 | 13 | 2023-08-02 | F090W/F277W | 418.7 x 4 | 1215 |
NGC 1448 | GO-1685 | 13 | 2023-08-02 | F150W/F277W | 526.1 x 4 | ” |
NGC 1448 | GO-1685 | 14 | 2023-18-02 | F090W/F277W | 418.7 x 4 | ” |
NGC 1448 | GO-1685 | 14 | 2023-18-02 | F150W/F277W | 526.1 x 4 | ” |
NGC 1559 | GO-1685 | 1 | 2023-06-30 | F090W/F277W | 418.7 x 4 | ” |
NGC 1559 | GO-1685 | 1 | 2023-06-30 | F150W/F277W | 526.1 x 4 | ” |
NGC 1559 | GO-1685 | 2 | 2023-07-15 | F090W/F277W | 418.7 x 4 | ” |
NGC 1559 | GO-1685 | 2 | 2023-07-15 | F150W/F277W | 526.1 x 4 | ” |
NGC 5584 | GO-1685 | 9 | 2023-01-30 | F090W/F277W | 418.7 x 4 | ” |
NGC 5584 | GO-1685 | 9 | 2023-01-30 | F150W/F277W | 526.1 x 4 | ” |
NGC 5584 | GO-1685 | 10 | 2023-02-21 | F090W/F277W | 418.7 x 4 | ” |
NGC 5584 | GO-1685 | 10 | 2023-02-21 | F150W/F277W | 526.1 x 4 | ” |
NGC 5643 | GO-1685 | 11 | 2023-07-07 | F090W/F277W | 311.4 x 4 | ” |
NGC 5643 | GO-1685 | 11 | 2023-07-07 | F150W/F277W | 418.7 x 4 | ” |
NGC 5643 | GO-1685 | 12 | 2023-07-22 | F090W/F277W | 311.4 x 4 | ” |
NGC 5643 | GO-1685 | 12 | 2023-07-22 | F150W/F277W | 418.7 x 4 | ” |
NGC 2525 | GO-2875 | 1 | 2023-04-23 | F090W/F277W | 472.4 x 3 | 1225 |
NGC 2525 | GO-2875 | 1 | 2023-04-23 | F150W/F277W | 526.1 x 3 | ” |
NGC 3370 | GO-2875 | 2 | 2023-06-28 | F090W/F277W | 472.4 x 4 | 1234 |
NGC 3370 | GO-2875 | 2 | 2023-06-28 | F150W/F277W | 526.1 x 4 | ” |
NGC 3447 | GO-2875 | 12 | 2023-05-24 | F090W/F277W | 472.4 x 4 | ” |
NGC 3447 | GO-2875 | 12 | 2023-05-24 | F150W/F277W | 622.7 x 4 | ” |
3 Photometry
We follow the general procedures outlined by previous works from our team for photometric reductions (Riess et al., 2024; Anand et al., 2024a). In detail, we perform PSF photometry333The photometry for this study is available publicly upon publication in a Zenodo repository, DOI: 10.5281/zenodo.13131990. using the DOLPHOT444http://americano.dolphinsim.com/dolphot/ software package (Dolphin, 2000, 2016), together with the latest (February 2024) version of the JWST/NIRCam module (Weisz et al., 2024). We use the stage 3 F150W i2d files as reference frames and perform photometry directly on the stage 2 cal images. We use the Vega-Vega system to remain consistent with earlier versions of DOLPHOT’s JWST module, which gives a 0.04 mag offset (in F090W) from the Vega-Sirius system that is currently default in DOLPHOT NIRCam module. We use the “-etctime” option, which allows the package to use (TMEASURE) for S/N calculations. We apply the following DOLPHOT quality cuts, similar to those from Weisz et al. (2024): crowd 0.5 (in both bands), sharp2 0.01 (in both bands), and type 2. In addition to these quality cuts, we also apply S/N cuts in both bands for each target. For NGC 1448, NGC 1559, NGC 3447, NGC 5584, and NGC 5643, we adopt a minimum threshold of S/N=5, whereas we choose a value of S/N=3 for NGC 2525 an NGC 3370 (due to the decreased relative depth of their photometry).
We also correct for foreground extinction using the Python package dustmaps
(Green, 2018; Green et al., 2019) and galaxy coordinates from the NASA/IPAC Extragalactic Distance Database (NED)555https://ned.ipac.caltech.edu/. We use the Fitzpatrick (1999) reddening law with = 3.1 and = 1.4156 and 0.6021 for F090W and F150W, respectively (Anand et al., 2024a).
4 Measurement
Measuring the TRGB magnitude involves locating the discontinuity or inflection point of the luminosity function consisting of red giant and asymptotic giant branch stars. This can be done many ways; for instance, using a Sobel filter (Lee et al., 1993; Hatt et al., 2017), or model-based methods, with either least-squares fitters (Wu et al., 2014; Crnojević et al., 2019) or maximum-likelihood estimation (Makarov et al., 2006; Li et al., 2022, 2023a). In this study, we measure the TRGB using a Sobel filter using the measurement routine without spatial clipping (which we instead perform using elliptical annuli) from the Comparative Analysis of TRGBs (CATs) team (Wu et al., 2023; Scolnic et al., 2023; Li et al., 2023b), which is publicly available on GitHub666https://github.com/JiaxiWu1018/Unsupervised-TRGB. The CATs team had optimized their measurement parameters (such as smoothing value, color cuts, etc.) by minimizing field-to-field dispersion in the TRGB measurement across multiple fields. At the moment, there are a limited number of JWST fields in each host galaxy, which prevents us from performing a similar optimization. For this study, we adopt a smoothing value of 0.05 mag and unweighted, or ‘simple’ weighting as named in the CATs algorithm (i.e. no further weighting of the output of the edge-detection measurement). We note that without the CATs optimizations in place, this methodology is equivalent to the TRGB measurements adopted by the Carnegie-Chicago Hubble Program (CCHP; Beaton et al., 2016; Freedman et al., 2019), but without the additional weighting of the Sobel filter output (which has been shown to cause biases in the outputs; Anderson et al. 2024; Anand et al. 2024a). We explore the effects of varying the level of smoothing of the luminosity function and the color range used to measure the luminosity function, 14 combinations in all, in the next section. TRGB uncertainties are estimated using bootstrap resampling of the luminosity function with 100 samples. For NGC 5584, using the smaller smoothing value of =0.05 mag produces a Sobel response that shows two peaks of nearly the same height. Following Anand et al. (2024a), if there are two Sobel peaks that are of equivalent height ( 2), we select the brighter peak.
Before measuring the TRGB, we perform spatial selections to limit contamination from younger stellar populations in these star-forming galaxies. The NIRCam regions used to measure the TRGB are highlighted in Fig. 1 in all cases (except NGC 3447, described below), only regions exterior to the mag/sq arcsec isophote (; de Vaucouleurs et al. 1991) are selected (Anand et al., 2022, 2024a; Hoyt et al., 2024). For NGC 1559 and NGC 5584, the parallel NIRCam module contains a suitable sampling of stars, although only half of the parallel module is used (the far half of each contain too few stars to perform precise empirical PSF adjustments and aperture corrections). The same is nearly true for NGC 5643, though we remove a small portion of each parallel module that is interior to . In NGC 1448, NGC 2525 and NGC 3370, a sizeable portion of the main module is outside of , which we include to increase the sampling of stars and for NGC 2525 and NGC 3370, the parallel modules contain too few stars to perform photometry on. Lastly, NGC 3447 is an actively interacting system, often described as two systems, N3447A and B. Here, we use two ellipses centered at RA, Dec = (10:53:23.998, +16:46:26.78) and (10:53:29:3934, +16:47:02.294), major axes of 140 and 50 arcsec, minor axes of 60 and 40 arcsec, and position angles of 80 and 115 degrees, respectively, following Mazzei et al. (2018). A color image of NGC 3370 is shown in Fig. 2, along with the same selection region adopted for the spatial cuts. It can be seen that the outer regions of the image used for our analysis are of low stellar density, where stars are easily separated.
The TRGB in F090W is expected to have minimal variation (0.02 mag) with color over a range of 1.15 F090W F150W 1.75 mag (Anand et al., 2024a). Here we adopt a color selection of 1.30 F090W F150W 1.75 mag for our baseline result (see Fig. 2 in Anand et al. 2024a), where the tighter range aims to limit the effects of a limited amount of supergiants which remain in our CMDs.
5 Distances
We use the TRGB measurements shown in Fig. 3 to measure distances to the host galaxies. We adopt a TRGB zero-point of = (stat) (sys) mag from Anand et al. (2024a) based on mag, which is the mean of the two simple Sobel measurements using a D and outer spatial selections (either being within 0.005 of this mean). This calibration uses the geometric maser distance to NGC 4258 (Reid et al., 2019), to calibrate the TRGB in the JWST F090W system and is stable for smoothing values mag to mag due to the absence of weighting (see Fig. 6 in Anand et al. 2024a). This absolute calibration is also highly consistent with the independently derived zero-point of = (stat) 0.043 (sys) mag from Newman et al. (2024b), after adjusting their value from the SiriusVega system to the VegaVega zeropoint system (M. Newman, priv. communication). We list and plot the distances to the host galaxies, compared to Cepheid-based distances from Riess et al. (2022). We compare to fit variant 10 from (Riess et al., 2022) for consistency in anchor selection; the distances produced by fit variant 10 are on average greater than those using all anchors (baseline) by 0.015 mag, consistent with their differences in of 72.5 km s-1 Mpc-1 for NGC 4258 and 73.0 km s-1 Mpc-1 using all anchors.
We find a weighted mean difference between the TRGB (baseline results with mag and smoothing of mag) and HST Cepheid distances of 0.01 0.04 (stat) 0.04 (sys) mag, which includes the systematic error due to measurement choice listed in Table 2, between the TRGB distance moduli measured here and Cepheid distance moduli from HST. We find no statistically significant difference between the two sets of distances measured between the same anchor and SN Ia hosts but with independent distance indicators. We have not attempted to account for any further population-based differences intrinsic to the TRGB between the field of NGC 4258 and the mean of the SN Ia hosts, but believe these are limited due to the limited metallicity and age differences expected in F090W and our restricted color window (Anand et al., 2024a).
Galaxy | TRGB | TRGB | TRGB | (TRGB) | (TRGB) | (Cepheid) | (Cepheid) | (var) | ||||
s=0.10 | s=0.10 | s=0.05 | s=0.05 | (s=0.10s=0.05) | s=0.10 | s=0.10 | s=0.05 | s=0.05 | ||||
NGC 1448 | 26.96 | 0.01 | 27.01 | 0.05 | 0.05 | 31.33 | 0.05 | 31.38 | 0.07 | 31.30 | 0.05 | 0.02 |
NGC 1559 | 27.12 | 0.01 | 27.13 | 0.01 | 0.01 | 31.49 | 0.05 | 31.50 | 0.05 | 31.50 | 0.07 | 0.01 |
NGC 2525 | 27.48 | 0.03 | 27.43 | 0.06 | 0.05 | 31.86 | 0.07 | 31.81 | 0.09 | 32.06 | 0.11 | 0.04 |
NGC 3370 | 27.83 | 0.02 | 27.84 | 0.08 | 0.01 | 32.20 | 0.06 | 32.21 | 0.09 | 32.13 | 0.06 | 0.02 |
NGC 3447 | 27.55 | 0.03 | 27.54 | 0.08 | 0.01 | 31.93 | 0.06 | 31.92 | 0.09 | 31.95 | 0.05 | 0.01 |
NGC 5584 | 27.58 | 0.06 | 27.43 | 0.10 | 0.15 | 31.95 | 0.08 | 31.8 | 0.11 | 31.77 | 0.06 | 0.07 |
NGC 5643 | 26.19 | 0.01 | 26.20 | 0.01 | 0.01 | 30.56 | 0.13 | 30.57 | 0.13 | 30.55 | 0.06 | 0.01 |
x
The choice of algorithm and selection parameters can impact final TRGB measurements. One compelling approach is to optimize the measurement parameters to minimize scatter in the field-to-field variations (see Wu et al. 2023; Scolnic et al. 2023; Li et al. 2023b), however, the limited number of JWST fields available to measure the TRGB within our host galaxies prevent us from performing such an optimization. To investigate the effects of how different parameter choices can impact our measurements, we explore a fourteen combinations of different luminosity function smoothing, color selection, and zero-point choice. We list the different measurement variants explored in this analysis, together with the weighted mean difference between the TRGB and HST Cepheid distances for each variant, in Table 4. We include a thirteenth variant using the zero-point from Newman et al. (2024b), adjusted for the VegaVega photometric system which has differences of 0.04 mag in F090W and 0.02 mag in F150W. We also add a variant where we use the median of all TRGB measurements (three smoothings, four color ranges) to compare with Cepheids. We add the standard deviation of the measured TRGBs across the smoothing and color range variants for each galaxy, excluding the variant using the zero-point from Newman et al. (2024b) and median TRGB, as a systematic error to our measurements, listed in Table 2. We find that for all variants, the mean differences between the TRGB and Cepheid distances across all galaxies for a given variant are consistent with zero to within their statistical uncertainties with a full range of 0.01 to 0.03 mag.
In Fig. 4, the error in the mean of 0.030.04 mag is a statistical uncertainty. There is also a systematic uncertainty in the mean TRGB and Cepheid comparison from the measurement uncertainty of the TRGB in NGC 4258 of mag (Anand et al., 2024a) which pertains to all calibrated TRGB distances and likewise for Cepheids in NGC 4258 of mag, or mag when they are combined. This systematic term due to the measurement of NGC 4258 for both methods is the dominant term in the comparison and may only be reduced in the future with the use of additional observations of NGC 4258. The maser distance uncertainty of mag should not be considered in this distance comparison as it is common to both the Cepheid and TRGB distances.
We can also compare the distances to NGC 1559 and NGC 5584 from Anand et al. (2024a), who used the same dataset as here. Anand et al. (2024a) found distances of = 31.49 0.07 mag ( = 0.01 mag) and 31.80 0.08 mag ( = 0.00 mag) to NGC 1559 and NGC 5584, respectively, when comparing their results to ours from the same smoothing scale (for NGC 1559, we now have combined both visits to obtain higher S/N photometry). The TRGB distance for NGC 1559 is also in agreement with that measured with Miras in Huang et al. (2020) of = 31.41 0.050 0.060 mag.
An unavoidable feature of the TRGB method of edge detection is a potential ambiguity when multiple Sobel peaks of similar height are present. For NGC 5584 there are two Sobel peaks separated by 0.20.3 mag. Because the peaks are similar in height, different measurement choices may cause changes in which peak is higher. To account for this possibility, we provide an additional individual distance uncertainty listed in Table 2 estimated from the variation of individual results of all the variants in Table 4. This uncertainty should be added for citing a single distance estimate for each host. A similar situation involving two peaks at similar heights, but at a smaller level, 0.05 mag, is seen for NGC 1448 and in the reverse direction for NGC 2525.
The Extragalactic Distance Database (EDD; Anand et al., 2021a) and CCHP (Freedman, 2021b) teams have measured HST TRGB distances to NGC 1448 and NGC 5643, which can provide an interesting comparison with the JWST distances found here. These distances also reflect differences in TRGB measurement methodology; the EDD team uses a model-based approach (Makarov et al., 2006; Wu et al., 2014), while the CCHP team uses a Sobel filter applied on a Gaussian-windowed, Locally Weighted Scatterplot Smoothing (GLOESS) smoothed luminosity function (Hatt et al., 2017). We list this comparison in Table 3. For NGC 1448, these teams independently found distance moduli of 31.38 mag (EDD) and 31.32 mag (CCHP), compared to 31.33 mag and 31.38 mag (different levels of smoothing) found here resulting in a standard deviation of 0.03 mag. For NGC 5643, these team found 30.47 mag (EDD) and 30.48 mag (CCHP), compared to 30.57 mag (which agrees well with Cepheids) found here resulting in a standard deviation of 0.04 mag.
Galaxy | Reference | (TRGB) |
NGC 1448 | Anand et al. (2021a) | 31.38 |
NGC 1448 | Freedman (2021b) | 31.32 |
NGC 1448 | Here | 31.33 (s=0.05), 31.38 (s=0.10) |
NGC 5643 | Anand et al. (2021a) | 30.47 |
NGC 5643 | Freedman (2021b) | 30.48 |
NGC 5643 | Here | 30.57 |
Smoothing [mag] | Color Cuts [mag] | [mag] | [mag] | Calibration | Weighted Scatter [mag] | |
0.05 | 1.15, 1.65 | 0.03 | 0.04 | Anand et al. (2024a) | 1.075 | 0.10 |
0.05 | 1.15, 1.75 | 0.03 | 0.04 | Anand et al. (2024a) | 1.468 | 0.12 |
0.05 | 1.30, 1.75 | 0.01 | 0.04 | Anand et al. (2024a) | 0.718 | 0.09 |
0.07 | 1.15, 1.65 | 0.02 | 0.04 | Anand et al. (2024a) | 0.919 | 0.09 |
0.07 | 1.15, 1.75 | 0.02 | 0.04 | Anand et al. (2024a) | 1.377 | 0.11 |
0.07 | 1.30, 1.75 | 0.02 | 0.04 | Anand et al. (2024a) | 1.142 | 0.10 |
0.10 | 1.15, 1.65 | 0.02 | 0.03 | Anand et al. (2024a) | 0.916 | 0.09 |
0.10 | 1.15, 1.75 | 0.02 | 0.03 | Anand et al. (2024a) | 1.081 | 0.09 |
0.10 | 1.30, 1.75 | 0.02 | 0.04 | Anand et al. (2024a) | 0.954 | 0.09 |
0.10 | 1.30, 1.75 | 0.01 | 0.04 | Newman et al. (2024b) | 0.946 | 0.08 |
0.15 | 1.15, 1.65 | 0.01 | 0.03 | Anand et al. (2024a) | 0.912 | 0.08 |
0.15 | 1.15, 1.75 | 0.02 | 0.03 | Anand et al. (2024a) | 0.878 | 0.08 |
0.15 | 1.30, 1.75 | 0.02 | 0.03 | Anand et al. (2024a) | 0.802 | 0.08 |
All Smoothings | All Color Ranges | 0.02 | 0.03 | Anand et al. (2024a) | 1.063 | 0.09 |
6 Discussion
We measure F090W TRGB-based distances to NGC 1448, NGC 1559, NGC 2525, NGC 3370, NGC 3447, NGC 5584, and NGC 5643 using JWST NIRCam observations and test their consistency with HST Cepheid-based distances to those same galaxies. We find excellent agreement between the two independent sets of distance measurements, which were measured between the same hosts and anchor but using data taken with different telescopes and distance indicators.
Examining the Hubble Tension necessitates close scrutiny of the consistency and systematics of distance indicators. Based on the measurements made in this study, we find no evidence of a bias or systematic difference between the Cepheid and TRGB methods that would cause a difference translatable to that can solve the Hubble Tension. This finding is in agreement with Freedman & Madore (2023), who find, “the excellent agreement between the published Cepheid distances in Riess et al. (2022) and TRGB distances in Freedman et al. (2021), which in the mean, agree to 0.007 mag”. In addition, Fig. 5 shows a comparison between Cepheid and TRGB distances, both measured with HST, using TRGB and Cepheid distances from Freedman (2021b) and Riess et al. (2022), respectively. The TRGB and Cepheid distances in Fig. 5 yield a mean difference of 0.00 0.02 (stat) 0.04 (sys) mag, also consistent with no difference even without additional consideration of populations differences in TRGB measurements (Anderson et al., 2024; Koblischke & Anderson, 2024). This comparison will benefit from more JWST TRGB measurements in the future (for instance, from GO-1995; Freedman et al. 2021). As the Hubble Tension corresponds to 5(73/67.5) 0.17 mag, the demonstrated consistency is meaningful and is inconsistent with providing evidence for a solution to the Hubble Tension at 3.
We do not yet report a value of from the JWST-calibrated SN subsamples because they are too small and the anchors too few to be competitive with the HST sample of 42 SN Ia and 4 anchors. The uncertainty in would be 2.5 times greater with 2-2.5 km/s/Mpc vs. 0.9 km/s/Mpc for HST, which would trivially remove the Hubble Tension simply by inflating the error in . Rather, the greater power of these presently small JWST samples comes from a direct comparison of what they can both measure in common, the distances between the same anchor and SN hosts.
It will be important to also scrutinize the difference in direct measurements that arise from different treatments of SNe Ia. Scolnic et al. (2023) find that the addition of peculiar flow corrections and cross-calibration of datasets which are included in the Pantheon and Pantheon+ analyses increase measurements each by 0.5 and 1.1 km/s/Mpc from the Freedman (2021a) analysis, bringing it closer in line with the value obtained from Cepheids. An under recognized source of variations in determinations of is also the make-up of the SN Ia calibration sample which can produce fluctuations of 1-2 km/s/Mpc due to small sample statistics.
Crosschecks between standard candles can be made more comprehensive with additional, independent standard candles such as Miras and carbon stars. The carbon star method, also called the J-region Asymptotic Giant Branch (JAGB), originated in the 1980s (Richer, 1981; Richer et al., 1984, 1985; Pritchet et al., 1987; Cook et al., 1986) and has since been revived and further pioneered in the 2000s (Battinelli & Demers, 2005; Ripoche et al., 2020; Madore & Freedman, 2020; Freedman & Madore, 2020; Parada et al., 2021, 2023; Zgirski et al., 2021; Lee et al., 2021b, a; Madore et al., 2022; Lee et al., 2022; Lee, 2023; Lee et al., 2024; Li et al., 2024). Although the JAGB method still requires further development and standardization (see discussions of asymmetric luminosity functions and metallicity effects in Parada et al. 2021, 2023; Li et al. 2024), the great luminosities of carbon stars in the near-infrared allow the JAGB to reach farther than the I and J-band TRGB, potentially allowing for a more comprehensive crosscheck of Cepheid variables via galaxy distances.
The TRGB measurements made here can benefit from further analysis, such as a future contrast ratio calibration (see CATs, Wu et al. (2023); Scolnic et al. (2023); Li et al. (2023b) or via modelling of the luminosity function (Makarov et al., 2006; Anand et al., 2019b, 2024a) instead of edge-detection algorithms (Anand et al., in prep). However, even without these additions, the consistency of the TRGB-based distance measurements with those of Cepheid-based distances measurement do not show evidence of a bias or systematic offset in Cepheid distances that would resolve the Hubble Tension. Similar analyses in the future will also benefit from further studies investigating the effects of red giant diversity to improve the accuracy of TRGB distances (such as in Anderson et al., 2024; Koblischke & Anderson, 2024). Future observations of TRGB fields with HST (for instance, from GO-17520; Breuval et al. 2023), JWST, and the Roman Space Telescope will provide more opportunities to scrutinize the second rung of the distance ladder by providing a more extensive set of observations that can be used to compile a more comprehensive comparison between these distance indicators. Specifically, wide-field JWST observations within the anchor galaxy NGC 4258 will help reduce systematic uncertainties with regards to the absolute calibration of the TRGB, and deep JWST observations of galaxies with the highest levels of mismatch between HST Cepheid and HST TRGB distances (e.g. NGC 3021, NGC 4038/9; Jang & Lee 2017b) will help elucidate any individual discrepancies between the two techniques. Lastly, a fully independent distance ladder from the traditional Cepheid+SNe Ia route is also under construction, via the usage of the TRGB combined with Surface Brightness Fluctuations as the final rung (Anand et al., 2024b).